OpenStudy (anonymous):

Can someone find an example of two divergent series whose double series is convergent? That is to say, \(\displaystyle\sum_{m=1}^\infty F(m,n)=\infty~\text{for all}~n\in\mathbb{N}\), and \(\displaystyle\sum_{n=1}^\infty F(m,n)=\infty~\text{for all}~m\in\mathbb{N}\), but \(\displaystyle\sum_{m=1}^\infty\sum_{n=1}^\infty F(m,n)<\infty\).

OpenStudy (anonymous):

\[\large\sum_{m=1}^\infty F(m,n)=\infty~\text{for all}~n\in\mathbb{N}\]\[\large\sum_{n=1}^\infty F(m,n)=\infty~\text{for all}~m\in\mathbb{N}\]\[\large\sum_{m=1}^\infty\sum_{n=1}^\infty F(m,n)<\infty\]I was thinking something along the lines of the classic "Gabriel's horn paradox", where the area under \({1\over x}\) from \(x=1\) to \(\infty\) is infinite, but the volume contained in the solid of revolution over the same interval is finite.

OpenStudy (anonymous):

That's not a very good example for this question though... :| Here, I'll draw a picture!

OpenStudy (anonymous):